OpenStudy (anonymous):

log[4](16)(√128) help me how to solve this ...

Parth (parthkohli):

Hey. Are you aware that \(\log(ab ) = \log( a) + \log (b)\)?

OpenStudy (anonymous):

nope i don't

Parth (parthkohli):

Hmm, weird. What identities do you know related to logarithms?

OpenStudy (anonymous):

like th quotient rule , power rule and the product rule?

Parth (parthkohli):

Wait, that sounds more like derivatives. Haha

OpenStudy (anonymous):

okay i have no idea .. Pls help me

OpenStudy (***[isuru]***):

there are some important things... \[\log_{a} a = 1\]\[\log \ b ^{a} = a \log \ b\]\[\log\ a + \log \ b = \log (ab)\]\[\log \ a - \log \ b = \log \ \frac{ a }{ b }\]

OpenStudy (***[isuru]***):

for the third and forth ones the base should be equal

OpenStudy (anonymous):

=108.9844999

OpenStudy (***[isuru]***):

now.. in ur problem to simplify it u can use the third one and write them separately as parthkoli done in the beginning

OpenStudy (anonymous):